OpenStudy (anonymous):

Find the missing endpoint and length of the line segment AB if the coordinates of A are (34,-22) and the coordinates of the midpoint are M(-2.5,10)? The coordinates of the missing endpoint of the line segment AB are B( , ) The length of the line segment AB (to the nearest tenth) = units

OpenStudy (anonymous):

let B=(x,y) (-2.5,10)=[(34+x)/2 , (-22+y)/2] therefore we have.... (34+x)/2 = -2.5 (-22+y)/2=10 34+x=-5 (-22+y)=20 x=-39 y=42 thus B's coordinate is (-39, 42)

OpenStudy (anonymous):

@mohammad.arqum Thx bro

OpenStudy (anonymous):

No problem...